Among 157 African-American men, the mean systolic blood pressure was 146 mm Hg with a standard deviation of 27. We wish to know if on the basis of these data, we may conclude that the mean systolic blood pressure for a population of African-American is greater than 140. Use α=0.01. 


1
Expert's answer
2021-12-16T16:07:33-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu\leq 140"

"H_1:\mu>140"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=157-1=156" degrees of freedom, and the critical value for a right-tailed test is "t_c =2.350489."

The rejection region for this right-tailed test is "R = \{t: t > 2.350489\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{146-140}{27/\sqrt{157}}\approx2.784436"

Since it is observed that "t = 2.784436>2.350489=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for right-tailed, "df=156, t=2.784436," is "p = 0.003013,"

and since "p = 0.003 013< 0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is greater than "140," at the "\alpha = 0.01" significance level.

Therefore, there is enough evidence to conclude that the mean systolic blood pressure for a population of African-American is greater than "140," at the "\alpha = 0.01" significance level.